OpenStudy (anonymous):

2(-4+7x)=-78 can you tell me what x is.

OpenStudy (saifoo.khan):

x is an English Alphabet.

OpenStudy (saifoo.khan):

2(-4+7x)=-78 -8 +14x = -78 14x = -78 + 8 14x = 70 x = 70/14 x = 10/2 x = 5